Super Carpeting Inc. just paid a dividend (D) of $3.36, and its dividend is expected to grow at a constant rate (g) of 4.90%1st drop down options: $47.89, $28.77, $45.71, $51.31

2nd drop down options: 5.13%, 7.02%, 7.70%, 7.35%

3rd drop down options: $47.89, $30.12, $50.20, $45.71

4th drop down options: 104.82%, 4.82%, 5.78%, 4.34%
Answer #1

rate positively ..

Ans a) Intrinsic value = Expectd dividend next year/(required rate - growth rate)
Expected dividend next year = 3.36*104.9%=         3.52
Therefore, Intrinsic value = 3.52/(12.25%-4.9%)
      47.89
Ans =       47.89
Ans b) Correct answer is option: The constant growth model implies that the dividend growth remains constant from now to infinity.
Ans c) Expected dividend yield = dividend next year/Price today
3.52/47.89
7.35%
Ans d) Price 1 year from today = Price today*(1+growth rate)
47.89*(1+4.9%)
50.2
Ans =       50.20
Ans e) Capital gain yield = (Price 1 year from now -Price today)/Price today
(50.2-47.89)/47.89
4.82%